alones (alones):

@johnweldon1993 Sorry :3

alones (alones):

Solve this quadratic equation using the quadratic formula. I NEED HELP WIHT THIS TOO \[2x^2 - 10x + 7 = 0\]

alones (alones):

@johnweldon1993 @johnweldon1993 @johnweldon1993 @johnweldon1993 @johnweldon1993 @johnweldon1993 @johnweldon1993 @johnweldon1993 @johnweldon1993 @johnweldon1993 @johnweldon1993 @johnweldon1993

alones (alones):

;)

OpenStudy (johnweldon1993):

Haha i get it i get it XD You tell me! we just did an extremely similar one...follow the same steps :) What are your 'a' 'b' and 'c'?

alones (alones):

POh shoot, yea yeaaaa I'll try

OpenStudy (johnweldon1993):

I''ll be here to check :D

alones (alones):

Okay be here

OpenStudy (johnweldon1993):

XD I wont leave! sheesh! :P

alones (alones):

Okay so i did a,b,c now i'm doing \[\sqrt{6^2 +4ac}\]

OpenStudy (johnweldon1993):

O.o are we doing the same problem?

OpenStudy (johnweldon1993):

\[\large 2x^2 - 10x + 7=0\] What are your 'a' 'b' and 'c' values? list em out :)

alones (alones):

LOL ..duhh..i jsut wrote a formula >.> here\[\sqrt{6x^2+4(10)(7)}\]

alones (alones):

AAA DAMN okay a= b=10 c=7

alones (alones):

Oh shoot a is 2

OpenStudy (johnweldon1993):

a = 2 right b = check this one again c = 7 yes

alones (alones):

Wait wouldn't it be \[\sqrt{(10)^2-4(2)(7)}\]

alones (alones):

b=10?

OpenStudy (johnweldon1993):

Nope, we are still working on the values for the moment :) a = 2 b = ????? c = 7 And your 'b' value...just think you have an equation that LOOKS like \(\large ax^2 + bx + 0\) *Notice how everythin is added!* Well in YOUR equation you have \(\large 2x^2 - 10x + 7 = 0\] Look how here we end up subtracting something! Meaning your b is actually -10! not 10!

OpenStudy (johnweldon1993):

I made a typo up there...should be \(\large ax^2 + bx + c = 0\)

OpenStudy (first_world_problems):

lol so you got this i guess

alones (alones):

Ughh tbh this is distracitng \(\large ax^2 + bx + c = 0\) Lol why it doesn't shows in Latex way >.> anyways, i dont' know what b is, i give up

OpenStudy (johnweldon1993):

So we have a=2 = -10 c = 7 Now plug those into \(\large x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\)

alones (alones):

wait it worked john, on my post not ur D:

alones (alones):

Ooooh it was -10 >.> anyways

OpenStudy (johnweldon1993):

mmhmm :) So now onto the quadratic formula with your known values :)

OpenStudy (first_world_problems):

wait why isn't it $$\\{a^2+b^2=c^2}$$

alones (alones):

I got \[\sqrt{-10+4(2)(7)}\]

OpenStudy (johnweldon1993):

@First_world_PROBLEMS that is the Pythagorean theorm, we are working with quadratics here :) @AloneS Not quite...also, dont focus on 1 part of the formula..you have the whole thing to work with \[\large x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] \[\large x = \frac{-(-10) \pm \sqrt{(-10)^2 - 4(2)(7)}}{2(2)}\]

OpenStudy (lncognlto):

@First_world_PROBLEMS That's the pythagorean theorem. Quite different to the quadratic formula.

OpenStudy (first_world_problems):

lol ooooooooooooooooooooohhhhhhhhhhhhhhhhhhhhhhhh....ok i knew that >.>

alones (alones):

I'm dead.Sucha forest... i got \[\frac{ 10\sqrt{44} }{ 4}\]

OpenStudy (first_world_problems):

ohh i can d that

alones (alones):

No i mean\[\frac{ 10\pm \sqrt{44} }{ 4 }\]

OpenStudy (johnweldon1993):

Remember you get 2 answers @AloneS type em both out for me :)

OpenStudy (usukidoll):

that square root can be split and simplified further

alones (alones):

OOOOOH AGAIN WEELL-DONE Fine fine fine i'll mke it \[\frac{ 10+\sqrt{44} }{ 4 },\frac{ 10-\sqrt{44} }{ 4 }\]

OpenStudy (johnweldon1993):

Perfect! Now! as @UsukiDoll has pointed out...that \(\large \sqrt{44}\) CAN be simplified further

alones (alones):

Oh nooo

OpenStudy (usukidoll):

yeeeeesssssssssss that's math life also after the split you can reduce the fraction too and split that up

alones (alones):

Usikii it is tortuing, literally i sit on each probelm 15 min each Oooh Oh lol nvm it can be broken to 11 and 4

OpenStudy (johnweldon1993):

Haha it's not bad at all Hint \(\large \sqrt{44} = \sqrt{4\times 11} = \sqrt{4}\sqrt{11}\) What do you notice? *think about what is under the square root

alones (alones):

oO.MG. 30 MINUTES I'M DOING THIS QUESTION D: Well-donneee i see they become negative

OpenStudy (usukidoll):

I just dealt with real analysis today. I'll teach you torture. Let f:D->R be continuous. Prove If D is bounded then f(D) is bounded

OpenStudy (first_world_problems):

5

OpenStudy (johnweldon1993):

Lol beat me to it! so perfect! now you would have \[\large \frac{10 + 2\sqrt{11}}{4}, \frac{10-2\sqrt{11}}{4}\] And you ca break that up and simplify further! :D haha

OpenStudy (johnweldon1993):

@UsukiDoll that looks like fun :D I just had to deal with my fluids final today >.< who wants to derive Navier Stokes for flow through a rough pipe :DD

alones (alones):

Oooh JOHN WHEN IS IT ENDING? I ca? john to what do i simplify it to?

OpenStudy (first_world_problems):

surely there's a shorter way

OpenStudy (usukidoll):

split the fractions and reduce

OpenStudy (usukidoll):

both are divisible by 2

alones (alones):

Uh-oh, 11 aND 2? usiki?

OpenStudy (johnweldon1993):

Well you have \[\large \frac{10 + 2\sqrt{11}}{4}, \frac{10 - 2\sqrt{11}}{4}\] Break that up into \[\large \frac{10}{4} + \frac{2\sqrt{11}}{4}, \frac{10}{4} - \frac{2\sqrt{11}}{4} \] Which gives you \[\large \frac{5}{2}+\frac{\sqrt{11}}{2},\frac{5}{2}-\frac{\sqrt{11}}{2}\] or \[\large \frac{5 + \sqrt{11}}{2}, \frac{5 - \sqrt{11}}{2}\]

OpenStudy (usukidoll):

^ #winning

alones (alones):

That's all O_O

OpenStudy (johnweldon1993):

Also @First_world_PROBLEMS it really is a short process, just a long post haha In general: in a nutshell 1) Find your coefficients, here a = 2, b = -10, c = 7 2) Plug them into the quadratic formula \(\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) 3) \(\large \frac{-(-10) \pm \sqrt{(-10)^2 - 4(2)(7)}}{2(2)}\) 4) \(\large \frac{10 \pm \sqrt{44}}{4}\) Then it's just simplifying

OpenStudy (usukidoll):

^ this user has a lot of patience... I don't lol

OpenStudy (johnweldon1993):

And @AloneS mmhmm that's all lol

OpenStudy (johnweldon1993):

Haha I'm studying for a final so ya know...procrastination at it's finest :D @UsukiDoll

OpenStudy (usukidoll):

my math professors curve so even if I eff up (which I try not to) I still pass so :P corruption at its finest.

OpenStudy (first_world_problems):

surely there's a shorter way

OpenStudy (johnweldon1993):

Ugh, I need the curve to be ever in my favor tomorrow >.< *When they force an ME major to take a PURELY EE based class* -sigh-

OpenStudy (usukidoll):

have you heard of Abstract Algebra?

OpenStudy (johnweldon1993):

Took it 2 semesters ago :D I blew through almost every math class offered here so far XD

alones (alones):

WaY TO GOOOOOO THANK YOU Ughhhh sucha torture

OpenStudy (usukidoll):

blew through?

OpenStudy (johnweldon1993):

@AloneS the question is...did you understand it? because if you ask another Quadratic problem I'm going to make you tell me every step >:D

OpenStudy (first_world_problems):

i got lost at the start.. idk why im still here

OpenStudy (usukidoll):

oh I got an even better one... if there's another quadratic problem, let's prove that f+g has a limit at x_0

OpenStudy (first_world_problems):

0_0

OpenStudy (johnweldon1993):

^lol :D

alones (alones):

You're here to spam @First_world_PROBLEMS ? O.MG. @johnweldon1993 I lost my breath when i read it :3 okay i'll return back to this post to look at the steps. I have the last question ;)

OpenStudy (first_world_problems):

NOPE BYE lol

OpenStudy (johnweldon1993):

Alright ladys and gents! I really SHOULD focus on studying XD that 8am final isn't coming any slower >.<

OpenStudy (usukidoll):

I had a 7:30 am exam...twice for foreign language

OpenStudy (johnweldon1993):

...that is disgusting on so many levels >.< lol

alones (alones):

Finally i got thrie the steps. Thanks everyone ;) (who helped._.)

OpenStudy (first_world_problems):

i wanted to but then i got lost and these redirects are sooo anoying